A cob Douglas production function for a firm is given as Q=4L ½K½. The firm has also established that wage rate and interest paid on capital are $3 and $5 respectively for a production period. The firm intents to spend $200 million for the period on production cost. Compute the levels of capital and labor that will maximize output. What is the maximum output?


1
Expert's answer
2021-03-09T15:20:10-0500

Solution:

Cost function: 200,000,000 = 3L + 5K

MPL = dQ/dL = 2L0.5-1K0.5 = 2L-0.5K0.5

MPK = dQ/dK = 0.5L0.5K0.5-1 = 0.5L0.5K-0.5

MPL/MPK = w/r

MPL/w = MPK /r

2L-0.5K0.5 / 3 = 0.5L0.5K-0.5 / 5

K = 15L

Plug into cost constraint:

200,000,000 = 3L + 5K

200,000,000 = 3L + (5x15L)

200,000,000 = 3L + 75L

200,000,000 = 78L

L = 2,564,102

K = 15L = (15x2564102) = 38,461,530

Maximum output = K + L = 38,461,530 + 2,564,102 = 41,025,632